# Learning in Markets with Heterogeneous Agents: Dynamics and Survival of Bayesian vs. No-Regret Learners

### v1

@@ 6.2 Bayesian Learning with Noisy Updates
Let $\epsilon>0$. The log-likelihood ratio under the noisy update rule takes the following form:
$\log\Big{(}\frac{P_{t}(\theta_{a})}{P_{t}(\theta_{b})}\Big{)}$ $=(1+\eta_{t})L(s_{t})+(1-\eta_{t})\log\Big{(}\frac{P_{t-1}(\theta_{a})}{P_{t-1}(\theta_{b})}\Big{)}$
$=(1+\eta_{t})\sum_{\tau=0}^{t-1}L(s_{t-\tau})\prod_{k=0}^{\tau}(1-\eta_{k})+\log\Big{(}\frac{P_{0}(\theta_{a})}{P_{0}(\theta_{b})}\Big{)}\prod_{\tau=0}^{t}(1-\eta_{\tau}),$ (26)
@@
$\log\Big{(}\frac{P_{0}(\theta_{a})}{P_{0}(\theta_{b})}\Big{)}\cdot(1-\eta)^{n_{+}(t)}(1+\eta)^{n_{-}(t)}=\log\Big{(}\frac{P_{0}(\theta_{a})}{P_{0}(\theta_{b})}\Big{)}\cdot\Big{[}\Big{(}\frac{1-\eta}{1+\eta}\Big{)}^{n_{+}(t)}\Big{]}\cdot(1+\eta)^{t}.$
@@
has probability less than $\delta$. Dividing by $(1+\eta)^{t}$ and taking the logarithm, we get
$n_{+}(t)\cdot\big{(}\log(1-\eta)-\log(1+\eta)\big{)}>\log(\epsilon)-t\log(1+\eta).$
Rearranging, we have
@@
Denote $c_{\epsilon}=\frac{\log(\epsilon)}{\log(1+\eta)-\log(1-\eta)}$, so
$n_{+}(t)<\frac{\log(1+\eta)}{\log(1+\eta)-\log(1-\eta)}\cdot t-c_{\epsilon}.$
Next, we observe that the coefficient of $t$ is strictly less than $1/2$. To see this, define
$c=\frac{1}{2}-\frac{\log(1+\eta)}{\log(1+\eta)-\log(1-\eta)}=-\frac{\log(1+\eta)+\log(1-\eta)}{2\big{(}\log(1+\eta)-\log(1-\eta)\big{)}}=\frac{-\log(1-\eta^{2})}{\text{Positive Number}}>0.$
Thus, we obtain that the event we want to bound is $n_{+}(t)<\Big{(}\frac{1}{2}-c\Big{)}t-c_{\epsilon}.$ Using Hoeffding’s inequality, using that $\operatorname{\mathbb{E}}[n_{+}]=\nicefrac{{t}}{{2}}$, we bound the probability of this event:
@@
is less than $\delta$, as required. Hence, the product approaches zero for large $t$: $\lim_{t\rightarrow\infty}\prod_{\tau=1}^{t}(1-\eta_{\tau})=0$ almost surely. ∎
@@
$\sum_{k=0}^{T}\binom{T}{k}p^{k}(1-\eta)^{k}\cdot(1-p)^{T-k}(1+\eta)^{T-k}=\big{(}p(1-\eta)+(1-p)(1+\eta)\big{)}^{T}=1\ \text{ (for $p=1/2$)}.$
The above claim implies that the second term in Equation (6.2) converges to zero almost surely. In other words, the learner forgets the prior, much like a perfect Bayesian learner (albeit at a somewhat faster rate in this case). For the first term of Equation (6.2)
@@
7 Simulations
Imperfect Bayesian Learning: Consider the following toy example. There is market with two assets where agent 1 is a Bayesian learner and agent 2 is a UCB learner, starting from equal wealth levels. [^10] We use UCB for simplicity. The main factor here is the different regret rates and not the particular algorithm. The distribution of states is $q=(0.7,0.3)$, i.e., state 1 occurs with probability $0.7$, and state 2 otherwise. The Bayesian learner has a prior consisting of three hypotheses: $(0.8,0.2)$, $(0.9,0.1)$, $(0.3,0.7)$, where the first hypothesis is closest to $q$ in relative entropy, but still has an error. The UCB learner considers a similar set of distributions as its action set: $(0.7,0.3)$, $(0.9,0.1)$, $(0.3,0.7)$, but the first of these being the correct distribution $q$. The example demonstrates the trade-off between short-term gains for the inaccurate Bayesian from converging quickly to a distribution close to the correct one, and the long-term dominance of the learner who converges to the correct distribution.
(a) Wealth Dynamics. The green oscillations show the step-by-step wealth of player 2. Solid lines show running window averages over the recent 10,000 steps in each time.
(b) Wealth distribution over time steps in the second half of the $10^{6}$ step simulation. The distribution shown in blue is the wealth of player 1 and in green is of player 2.
(c) Wealth distribution over time steps in the 10,000 initial steps of the dynamic. The distribution shown in blue is the wealth of player 1 and in green is of player 2.
Figure 1: Example of wealth dynamics in competition between an inaccurate Bayesian with an error in its support learner and regret minimizing agent who converges to the correct state distribution $q$.
Dynamics with a Constant Regret Difference: Next, we consider an example of two Bayesian learners, both with priors including the correct distribution of states. By Theorem 5.4, their regret levels can differ only by an additive constant, as both are perfect Bayesians with constant regret. Consequently, these agents are bound to coexist in the market. The example demonstrates that while the quality of the priors—provided they include a correct model—does not affect survival, the initial learning phase, which depends on the prior, does influence the steady-state partition of wealth between the agents. In this example, as in the previous one, there are two states, with the underlying distribution given by $q=(0.7,0.3)$. Agent 1 has a prior that includes the correct distribution and two other models: $(0.8,0.2),(0.7,0.3),(0.6,0.4)$. Agent 2’s prior includes one additional model: $(0.8,0.2),(0.7,0.3),(0.6,0.4),(0.5,0.5)$. Both agents start with a uniform prior over the options they consider. Thus, agent 2 places smaller weight initially on the correct model than does agent 1. Figure 2(a) depicts the wealth dynamics of these learners, starting form equal wealth levels.
(a) Two Bayesian learners.
(b) Two no-regret learners.
(c) Bayesian vs. no-regret learner.
Acknowledgments
Éva Tardos was supported in part by AFOSR grant FA9550-23-1-0410, AFOSR grant FA9550-231-0068, and ONR MURI grant N000142412742.

### v2

@@ Appendix A Proofs
(Theorem 3.1) This result follows from the analysis leading to Equation (16). For the first part of the theorem, we have $R^{n}(T)-R^{m}(T)=-\log(r^{nm})-\log(r_{0}^{nm})=\log(r^{mn})+const$. If the left-hand side diverges to +$\infty$+ ∞, then $r^{mn}=w^{m}/w^{n}$ diverges as well, which is only possible if $w^{n}\rightarrow 0$. For the second part, clearly this is a necessary condition due to the first part of the theorem: if an agent vanishes, it does not survive. To see that this condition is sufficient, if there exists a constant $c>0$ such that $R^{n}(T)-R^{m}(T)<c$ at all times and for all $m\neq n$, then the wealth ratio $r^{mn}$ with any other player is bounded, and thus $w^{n}$ is bounded from below. ∎
Proof.
(Theorem 5.2) Consider an agent indexed by $q$ using the state distribution $q$ as its investment strategy; i.e, $\alpha^{q}_{t}=q$ for all $t$, and we denote this constant sequence by $q_{1:T}$. Agent $q$ competes with a magic agent. The regret of the agent playing strategy $q$ is denoted $R^{T}(q_{1:T})$, and denote its expectation by $R=\operatorname{\mathbb{E}}[R^{T}(q_{1:T})]$. On the one hand, using the fact that $\hat{q}$ has zero regret by definition, from Equation 16 we have
$R^{T}(q_{1:T})=\log(r^{\hat{q}q}_{T})-\log(r^{\hat{q}q}_{0}).$
On the other hand, taking expectation of Equation 6, we have
$\operatorname{\mathbb{E}}\big{[}$ $\log(r^{\hat{q}q}_{T})\big{]}$
$=\operatorname{\mathbb{E}}\big{[}\sum_{s}\sum_{t}\mathbf{1}_{s_{t}=s}\log(\frac{\alpha^{\hat{q}}_{s}}{q_{s}})\big{]}+\log(r^{\hat{q}q}_{0})$
$=T\operatorname{\mathbb{E}}\big{[}\sum_{s}\frac{n^{s}_{T}}{T}\log(\frac{n^{s}_{T}/T}{q_{s}})\big{]}+\log(r^{\hat{q}q}_{0})$
$=T\cdot\operatorname{\mathbb{E}}[I_{\hat{q}}(q)]+\log(r^{\hat{q}q}_{0}).$
where $n^{s}_{T}$ is the empirical count of state $s$ until time $T$. And thus, we get $R=T\cdot\operatorname{\mathbb{E}}\big{[}I_{\hat{q}}(q)\big{]}$. That is, the expected regret equals the relative entropy (a.k.a. KL-divergence) between an empirical distribution and the state distribution scaled by $T$. The question of survival of this agent now hinges on how does $I_{\hat{q}}(q)$ depend on time. Here, we use a result from information theory, showing that the expected KL-divergence of the empirical distribution decreases as the alphabet size divided by the sample size [^10] We are interested in the expectation. [74] also provide concentration bounds for this result. (see, e.g., Lemma 13.2 in [79]). In our case, $\operatorname{\mathbb{E}}[I_{\hat{q}}(q)]\equiv\operatorname{\mathbb{E}}[D_{KL}(\hat{q}||q)]=\frac{S-1}{2T}+o(\frac{1}{T})$. Thus, the expected regret $R$ is constant, depending only on the number of states. ∎
Proof.
(Theorem 5.4) By Theorem 5.2, an agent $n$ using $\alpha^{n}_{t}=q$ for all $t$ incurs constant expected regret. By Proposition 4.2, a Bayesian learner whose support includes the distribution $q$ almost surely survives against such a competitor. Next, by Theorem 3.1 if some agent survives against a competitor, then the difference between their regret levels must be bounded by a constant at all times. Therefore, the Bayesian competing against the state distribution $q$ almost surely incurs constant regret. Having established the first point, the second follows directly from Theorem 3.1: Since the Bayesian has constant regret, any competitor who survives must also have constant regret. ∎
Proof.
(Theorem 6.1) Let $n$ be the index of a Bayesian agent with $q$ not within its support, and let $q^{\prime}\neq q$ be the strategy within the support that is closest to $q$ in relative entropy. By Proposition 4.1, the strategy used by the Bayesian agent, $\alpha^{n}_{t}$, converges almost surely to $q^{\prime}$. Hence, intuitively, for sufficiently large $t$, the strategies remain bounded away from $q$, leading to linear regret from that point onward. In more detail, define $I_{0}=\frac{1}{2}I_{q}(q^{\prime})>0$. In every sample path $i$ (i.e., an infinite sequence, indexed by $i$, of state realizations and the ensuing sequence of $\alpha^{n}_{t}$), except on a set of measure zero, there exists a time $T_{i}$ such that for all $t\geq T_{i}$, we have $I_{q}(\alpha^{n}_{t})>I_{0}$. The regret accumulated up to time $T_{i}$ is constant (possibly positive or negative). The expected regret difference relative to strategy $q$ at $t>T_{i}$ is $\operatorname{\mathbb{E}}[U^{t}(q)-U^{t}(\alpha^{n}_{t})]\geq I_{0}$, and so the expected regret relative to $q$ until time $T$ is at least $I_{0}\cdot(T-T_{i})+const$. By Theorem 5.2, the strategy $q$ incurs only constant regret, so the regret of the Bayesian agent converging to $q^{\prime}$ grows linearly in $T$. Now, consider a no-regret learner. Such a learner has sublinear worst-case regret and thus sublinear expected regret. By Theorem 3.1, the imperfect Bayesian agent who has linear regret vanishes. ∎
Proof.
(Theorem 6.5) Let $\epsilon>0$. The log-likelihood ratio under the noisy update rule takes the following form:
$\log\Big{(}\frac{P_{t}(\theta_{a})}{P_{t}(\theta_{b})}\Big{)}$ $=(1+\eta_{t})L(s_{t})$
$+(1-\eta_{t})\log\Big{(}\frac{P_{t-1}(\theta_{a})}{P_{t-1}(\theta_{b})}\Big{)}$
$=(1+\eta_{t})\sum_{\tau=0}^{t-1}L(s_{t-\tau})\prod_{k=0}^{\tau}(1-\eta_{k})$ (27)
$+\log\Big{(}\frac{P_{0}(\theta_{a})}{P_{0}(\theta_{b})}\Big{)}\prod_{\tau=0}^{t}(1-\eta_{\tau}),$
@@
$\log\Big{(}\frac{P_{0}(\theta_{a})}{P_{0}(\theta_{b})}\Big{)}\cdot(1-\eta)^{n_{+}(t)}(1+\eta)^{n_{-}(t)}$
$=\log\Big{(}\frac{P_{0}(\theta_{a})}{P_{0}(\theta_{b})}\Big{)}\cdot\Big{[}\Big{(}\frac{1-\eta}{1+\eta}\Big{)}^{n_{+}(t)}\Big{]}\cdot(1+\eta)^{t}.$
@@
has probability less than $\delta$. Dividing by $(1+\eta)^{t}$ and taking the logarithm, we get $n_{+}(t)\cdot\big{(}\log(1-\eta)-\log(1+\eta)\big{)}>\log(\epsilon)-t\log(1+\eta).$. Rearranging, we have
@@
Denote $c_{\epsilon}=\frac{\log(\epsilon)}{\log(1+\eta)-\log(1-\eta)}$, so $n_{+}(t)<\frac{\log(1+\eta)}{\log(1+\eta)-\log(1-\eta)}\cdot t-c_{\epsilon}.$ Next, we observe that the coefficient of $t$ is strictly less than $1/2$. To see this, define
$c$ $=\frac{1}{2}-\frac{\log(1+\eta)}{\log(1+\eta)-\log(1-\eta)}$
$=-\frac{\log(1+\eta)+\log(1-\eta)}{2\big{(}\log(1+\eta)-\log(1-\eta)\big{)}}$
$=\frac{-\log(1-\eta^{2})}{\text{Positive Number}}>0.$
Thus, we obtain that the event we wish to bound is $n_{+}(t)<\Big{(}\frac{1}{2}-c\Big{)}t-c_{\epsilon}.$ Using Hoeffding’s inequality, using that $\operatorname{\mathbb{E}}[n_{+}]=\nicefrac{{t}}{{2}}$, we bound the probability of this event:
@@
is less than $\delta$, as required. Hence, the product approaches zero for large $t$: $\lim_{t\rightarrow\infty}\prod_{\tau=1}^{t}(1-\eta_{\tau})=0$ almost surely, proving the claim. ∎
@@
$\sum_{k=0}^{T}\binom{T}{k}p^{k}(1-\eta)^{k}\cdot(1-p)^{T-k}(1+\eta)^{T-k}=\big{(}p(1-\eta)+(1-p)(1+\eta)\big{)}^{T}=1\ \text{ (since $p=1/2$)}.$
The above claim implies that the second term in Equation (A) converges to zero almost surely. In other words, the learner forgets the prior, much like a perfect Bayesian learner (albeit at a somewhat faster rate in this case). For the first term of Equation (A)
@@
Proof.
(Lemma 5.5) The result follows from the definition of regret as a utility difference, as discussed in Section 3. Specifically, the regret of strategy sequence $\alpha^{n}_{1:T}$ is $R^{T}(\alpha^{n}_{1:T})=U(\alpha^{\hat{q}}_{1:T})-U(\alpha^{n}_{1:T})$, also, $R=\operatorname{\mathbb{E}}[U(\alpha^{\hat{q}}_{1:T})-U(q_{1:T})]$, and denote the regret of $\alpha^{n}_{1:T}$ with respect to the benchmark $q_{1:T}$ as $R^{\alpha}_{q}=U(q_{1:T})-U(\alpha^{n}_{1:T})$. We have $\operatorname{\mathbb{E}}[R^{T}(\alpha^{n}_{1:T})]=\operatorname{\mathbb{E}}[R^{\alpha}_{q}]+R=\operatorname{\mathbb{E}}\Big{[}T\sum_{s}\frac{n^{s}_{T}}{T}\log(\frac{q_{s}}{\alpha^{n}_{s}})\Big{]}+R=T\cdot I_{q}(\alpha^{n})+R$. ∎
Proof.
(Proposition 7.1) Consider a market process with two states and $T$ time steps. Let $11/2<p<$ 1, and define $q_{1}=(p,1-p)$, $q_{2}=(1-p,p)$, and $Q=\{q_{1},q_{2}\}$. Suppose that for the first $T_{1}$ steps the data-generating distribution is $q_{1}$, and for the remaining $T_{2}=T-T_{1}$ steps it is $q_{2}$. At any time $t\leq T_{1}$, taking expectation in Equation (19), we have for a Bayesian learner starting from a uniform prior:
$\operatorname{\mathbb{E}}[\log\frac{\lambda^{1}_{t}}{\lambda^{2}_{t}}]=\operatorname{\mathbb{E}}[n^{1}_{t}]\cdot\log\frac{p}{1-p}+\operatorname{\mathbb{E}}[n^{2}_{t}]\cdot\log\frac{1-p}{p},$
where, recall, $n^{s}_{t}$ denotes the number of times state $s$ has occurred up to time $t$. At time $T_{1}$ we get:
$\operatorname{\mathbb{E}}[\log\frac{\lambda^{1}_{T_{1}}}{\lambda^{2}_{T_{1}}}]=pT_{1}\cdot\log\frac{p}{1-p}+(1-p)T_{1}\cdot\log\frac{1-p}{p}.$
After the shift, at $t=T_{1}+\tau$, using the same formula and the learner’s state at time $T_{1}$, we have:
$\operatorname{\mathbb{E}}[\log\frac{\lambda^{1}_{t}}{\lambda^{2}_{t}}]=\Big{(}pT_{1}\cdot\log\frac{p}{1-p}+(1-p)T_{1}\cdot\log\frac{1-p}{p}\Big{)}+\Big{(}(1-p)\tau\cdot\log\frac{1-p}{p}+p\tau\cdot\log\frac{p}{1-p}\Big{)}.$
Grouping terms, we get:
$\operatorname{\mathbb{E}}[\log\frac{\lambda^{1}_{t}}{\lambda^{2}_{t}}]=p(T_{1}+\tau)\cdot\log\frac{p}{1-p}+(1-p)(T_{1}+\tau)\cdot\log\frac{1-p}{p}.$
Intuitively, for any $\tau<T_{1}$, we have $\operatorname{\mathbb{E}}[n^{1}_{T_{1}+\tau}]>\operatorname{\mathbb{E}}[n^{2}_{T_{1}+\tau}]$, and so $\lambda^{1}_{t}>\lambda^{2}_{t}$ during steps $t=T_{1},\dots,T_{1}+\tau$. This implies that during the second phase, the learner still assigns probability greater than $1/2$ to state 1, despite state 1 occurring with probability less than $1/2$. This discrepancy leads to regret proportional to $\tau$. Concretely, take $T_{1}=2T/3$, $p=3/4$, and $\tau=T/3$. Then during steps $t=T_{1}+1,\dots,T$ we have:
$\operatorname{\mathbb{E}}[\log\frac{\lambda^{1}_{t}}{\lambda^{2}_{t}}]>\operatorname{\mathbb{E}}[\log\frac{\lambda^{1}_{T}}{\lambda^{2}_{T}}]=\frac{3T}{4}\cdot\log 3-\frac{T}{4}\cdot\log 3=\frac{\log 3}{2}T>\frac{1}{2}T.$
So the empirical probability the learner assigns to state 1 is at least $1/2$, whereas the true probability during the last $T/3$ steps is $1/4$. Therefore, similarly to the proof of Theorem 6.1, since there is a gap between the assigned and true probabilities, the regret accumulated during these steps after $32T/$ 3 is linear; in expectation it is at least the regret of the strategy $(\tfrac{1}{2},\tfrac{1}{2})$, which is
$\frac{T}{3}\cdot I_{q_{2}}\left(\tfrac{1}{2},\tfrac{1}{2}\right)=\frac{3\log 3+4\log 2}{4}\cdot\frac{T}{3}>\frac{T}{23}.$
During the first stage, until $32T/$ 3, the learner incurs a constant regret, so the total regret is linear in $T$. ∎
Proof.
(Theorem 7.2) Let us consider a phase of the process in which the data generating distribution is $q$. For notational convenience, we overload the symbol $q$ to also denote the index of this distribution, so that $\lambda^{q}$ denotes the weight the learner assigns to this correct model. By the update rule (26), for any $k\neq q$:
$\log\frac{\lambda^{k}_{t+1}}{\lambda^{q}_{t+1}}=\log\frac{\tilde{\lambda}^{k}_{t+1}+\epsilon_{t}}{\tilde{\lambda}^{q}_{t+1}+\epsilon_{t}}=\log\frac{\tilde{\lambda}^{k}_{t+1}}{\tilde{\lambda}^{q}_{t+1}}+\log\big{(}1+\tfrac{\epsilon_{t}}{\tilde{\lambda}^{k}_{t+1}}\big{)}-\log\big{(}1+\tfrac{\epsilon_{t}}{\tilde{\lambda}^{q}_{t+1}}\big{)}.$
Since for any $x>0$ it holds that $0<\log(1+x)\leq x$, we have
$\log\frac{\lambda^{k}_{t+1}}{\lambda^{q}_{t+1}}\leq\log\frac{\tilde{\lambda}^{k}_{t+1}}{\tilde{\lambda}^{q}_{t+1}}+\frac{\epsilon_{t}}{\tilde{\lambda}^{k}_{t+1}}.$
Applying the Bayesian update rule for the log-odds and taking expectations, we get
$\operatorname{\mathbb{E}}\Big{[}\log\frac{\lambda^{k}_{t+1}}{\lambda^{q}_{t+1}}\Big{]}\leq\log\frac{\lambda^{k}_{t}}{\lambda^{q}_{t}}-I_{q}(\theta^{k})+\operatorname{\mathbb{E}}\Big{[}\frac{\epsilon_{t}}{\tilde{\lambda}^{k}_{t+1}}\Big{]}.$
Now consider two cases. First, suppose first that $\tilde{\lambda}^{k}_{t+1}<t^{-\tfrac{3}{2}}$. For $t$ large enough (e.g., $t>6$), we have
$\lambda^{k}_{t+1}<\tilde{\lambda}^{k}_{t+1}+\epsilon_{t}\leq t^{-\tfrac{3}{2}}+t^{-2}<t^{-\tfrac{4}{3}}.$
Intuitively, once $\tilde{\lambda}^{k}_{t+1}$ becomes small, it stays small. In the second case, $\tilde{\lambda}^{k}_{t+1}\geq t^{-\tfrac{3}{2}}$. Then, we have
$\operatorname{\mathbb{E}}\Big{[}\log\frac{\lambda^{k}_{t+1}}{\lambda^{q}_{t+1}}\Big{]}\leq\log\frac{\lambda^{k}_{t}}{\lambda^{q}_{t}}-I_{q}(\theta^{k})+O\big{(}t^{-\frac{1}{2}}\big{)}.$ (28)
Since $I_{q}(\theta^{k})$ is a positive constant for any $\theta^{k}\neq q$, we see that for large enough $t$, the log-odds of $k\neq q$ relative to $q$ decrease at a linear rate, implying that the posterior distribution approaches exponentially fast the true model $q$. Thus, after a number of steps depending logarithmically on the value of $\lambda^{q}$ at the start of the phase, we reach—with high probability over the randomness in the data-generating process—the first case where $\tilde{\lambda}^{k}_{t+1}<t^{-\tfrac{3}{2}}$.
Constant regret for a stationary process: Given that $\lambda_{0}^{q}$ is a constant, after some constant time $T_{0}$ (independent of $T$), it holds with high probability that for all $t>T_{0}$ and all $k\neq q$, we have $\lambda^{k}_{t}<t^{-\tfrac{4}{3}}$. Using this, we can bound the regret of player $n$ under our strategy. Let $I$ denote the relative entropy of the worst model in the prior: $I=\max_{k}I_{q}(\theta^{k})$. With high probability,
$R^{T}(\alpha_{1:T}^{n})-R\leq 2\sum_{t=1}^{T}I_{q}(\alpha^{n}_{t})\leq 2\sum_{t=1}^{T}I\cdot t^{-4/3}<2\sum_{t=1}^{\infty}I\cdot t^{-4/3}=2I\cdot\zeta(4/3)\approx 2I\cdot 3.601.$
Where $\zeta(x)=\sum_{n=1}^{\infty}\tfrac{1}{n^{x}}$ denotes the Riemann zeta function which has finite value for $x>1$.
Logarithmic regret after a distribution shift: The above analysis shows one side of the regularization method: when the weights on models $k\neq q$ are large, they decay exponentially fast, similar to the standard Bayesian update.
The other side is what happens after a distribution shift. The key point is that while the weights on incorrect models become small enough to ensure constant regret, the regularization prevents them from becoming exponentially small. Suppose that after some history of length $T_{1}$, the stochastic process shifts to a new distribution, which we denote by $q^{\prime}$. A lower bound on $\lambda^{q^{\prime}}_{T_{1}+1}$ is the amount of regularization added to it in the previous step, namely,
$\lambda^{q^{\prime}}_{T_{1}+1}\geq\frac{\epsilon_{T_{1}}}{1+K\epsilon_{T_{1}}}\geq\frac{1}{K{T_{1}}^{2}}.$
The last inequality holds since the shift occurs at $t>1$.
So at time $T_{1}+1$, we can think about the situation as a new learning process with a prior weight of at least $\lambda^{q^{\prime}}_{T_{1}+1}\geq\frac{1}{K{T_{1}}^{2}}$. Because of exponential convergence, the time to “forget” this prior is logarithmic in $T_{1}$. Specifically, the exponent for model $k$ is $I_{q^{\prime}}(\theta^{k})$. Let $I$ now be the minimal entropy over models $k\neq q^{\prime}$. The expected time $\tau$ it takes to reach $\lambda_{t}^{k}\leq\lambda_{t}^{q^{\prime}}$ for all $k\neq q^{\prime}$ and eliminate this prior is $\tau=\tfrac{2}{I}\log T_{1}+\text{constant}$. After this time, we already know that the regret incurred until the next distribution shift, or reaching the time horizon $T$, is bounded by a constant independent of $T$. ∎